OpenStudy (anonymous):

6 over 7 divided by 3=

OpenStudy (anonymous):

What you are essentially doing is: \[\frac{6}{7} \times \frac{1}{3}\] Do you know how to do this?

OpenStudy (anonymous):

Not really

OpenStudy (anonymous):

To multiply 2 fractions together all you do is multiply the numerators together and the denominators together. So what you ahve to do is: \[\frac{6\times1}{7\times3}\] Understand?

OpenStudy (anonymous):

Yes thanks

OpenStudy (anonymous):

Excellent :) What do you get?

OpenStudy (anonymous):

6/21

OpenStudy (anonymous):

Excellent, can you simplify this fraction at all?

OpenStudy (anonymous):

3/21?

OpenStudy (anonymous):

Not quite, remember what you do to the top you have to do to the bottom as well if you want to keep the fraction the same. Hint: Both the numerator and denominator are multiples of 3.

OpenStudy (anonymous):

1/7?

OpenStudy (anonymous):

That would be right for 3/21, you looked at the wrong fraction. What we have is 6/21 :) Try again.

OpenStudy (anonymous):

3/7

OpenStudy (anonymous):

?

OpenStudy (anonymous):

No, we divide the top and bottom by 3. So we have: \[\frac{6}{21}=\frac{6\div3}{21\div3}=\frac{2}{7}\]

OpenStudy (anonymous):

Oh ok thanks